Exercise 10-4 Direct Labor and Variable Manufacturing Overhead Variances [LO10-2, LO10-3]

Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows:

Standard
Hours
Standard Rate
per Hour
Standard
Cost
24 minutes $5.60 $2.24

During August, 8,180 hours of direct labor time were needed to make 19,200 units of the Jogging Mate. The direct labor cost totaled $44,172 for the month.

Required:

1. What is the standard labor-hours allowed (SH) to makes 19,200 Jogging Mates?

2. What is the standard labor cost allowed (SH × SR) to make 19,200 Jogging Mates?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

5. The budgeted variable manufacturing overhead rate is $4.70 per direct labor-hour. During August, the company incurred $45,808 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month.

(For requirements 3 through 5, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1
1 Standard labor-hours allowed 7680 =19200*24/60
2 Standard labor cost allowed 43008 =7680*5.60
3 Labor spending variance 1164 U =44172-43008
4 Labor rate variance 1636 F =44172-(8180*5.6)
Labor efficiency variance 2800 U =5.6*(8180-7680)
5 Variable overhead rate variance 7362 U =45808-(8180*4.7)
Variable overhead efficiency variance 2350 U =4.7*(8180-7680)